What is the kinetic energy of an electron having a rest mass of 9.1 x 10^-31 kg traveling at 65% speed of light?

1
Expert's answer
2022-02-08T15:10:02-0500

"W=E-E_0=mc^2-m_0c^2=\Delta mc^2=(\frac 1{\sqrt{1-\frac{v^2}{c^2}}}-1)m_0c^2,"

"W=0.16~MeV."